OpenStudy (anonymous):

(2/3)+(1/2) Would the answer be (7/6)?

OpenStudy (raffle_snaffle):

Find a common denominator

OpenStudy (raffle_snaffle):

What is your common denominator?

OpenStudy (anonymous):

So I times (2/3) by 2 and (1/2) by 3

OpenStudy (anonymous):

Lol.

OpenStudy (raffle_snaffle):

yep, that is right.

OpenStudy (anonymous):

The best way to put it is: 1 1/6

OpenStudy (anonymous):

Only if the question is asking for a certain type of fraction.

OpenStudy (anonymous):

Thanks

OpenStudy (isaiah.feynman):

@staldk3 \[\frac{ a }{b} \pm \frac{ c }{ d} = \frac{ ad \pm bc}{ bd }\] You can also use this simple rule when the denominators of the two fractions are NOT the same.

OpenStudy (anonymous):

Attack on titan?

OpenStudy (isaiah.feynman):

@pak_asfand what are you talking about?

OpenStudy (anonymous):

It's his/her profile picture...its an anime attack on titan

OpenStudy (anonymous):

It's a good show.

OpenStudy (anonymous):

Lol. I'm not a fan of Hero o:

OpenStudy (anonymous):

Well now I am :P